# Reinforce Lifelong Interaction Value of User-Author Pairs for Large-Scale Recommendation Systems

### v1

@@ Abstract.
Recommendation systems (RS) help users find interested content and connect authors with their target audience. Most research in RS tends to focus either on predicting users’ immediate feedback (like click-through rate) accurately or improving users’ long-term engagement. However, they ignore the influence for authors and the lifelong interaction value (LIV) of user-author pairs, which is particularly crucial for improving the prosperity of social community in short-video platforms. Currently, reinforcement learning (RL) can optimize long-term benefits and has been widely applied in RS. In this paper, we introduce RL to Reinforce Lifelong Interaction Value of User-Author pairs (RLIV-UA) based on each interaction of UA pairs. To address the long intervals between UA interactions and the large scale of the UA space, we propose a novel Sparse Cross-Request Interaction Markov Decision Process (SCRI-MDP) and introduce an Adjacent State Approximation (ASA) method to construct RL training samples. Additionally, we introduce Multi-Task Critic Learning (MTCL) to capture the progressive nature of UA interactions (click → follow → gift), where denser interaction signals are leveraged to compensate for the learning of sparse labels. Finally, an auxiliary supervised learning task is designed to enhance the convergence of the RLIV-UA model. In offline experiments and online A/B tests, the RLIV-UA model achieves both higher user satisfaction and higher platform profits than compared methods.
@@ 1. Introduction
The recommendation system (RS) aims to explore users’ interested content, while helping content authors reach potential target users to accumulate fans and obtain profits (Zangerle and Bauer, 2022; Jesse and Jannach, 2021; Afsar et al., 2022). By promoting repeated interactions from both users and authors, RS helps build a more active ecosystem, leading to increased platform traffic and commercial returns (Azaria et al., 2013; Dias et al., 2008; Lu et al., 2014; Ras et al., 2017; Jannach and Adomavicius, 2017; Jannach and Jugovac, 2019; Jannach and Bauer, 2020).
@@
In the current field of RS, some research focuses on improving the accuracy of user immediate feedback prediction (Linden et al., 2003; He et al., 2014; Covington et al., 2016; Zhou et al., 2018; Gu et al., 2020), i.e. click-through rate (CTR), at each request by deep neural network (DNN) model, as shown in Fig. 1. Other studies concentrate on optimizing long-term engagement from the user’s perspective (Zhang et al., 2022; Xue et al., 2023; Chen et al., 2021; Wang et al., 2022; Cai et al., 2023; Zheng et al., 2018; Zou et al., 2019; Wang, 2021). Specifically, they utilize reinforcement learning (RL) to dynamically optimize session-level long-term cumulative rewards (Sutton, 2018; Van Hasselt et al., 2016; Fujimoto and Gu, 2021). However, the aforementioned research ignores the recommendation impact on authors and the Lifelong Interaction Value (LIV) of user-author (UA) pairs. Therefore, it is impossible to characterize the progressive changes of the lifelong UA relationship, which is essential for improving the stickiness (retention) of UA pairs and the prosperity of social community in short-video platforms.
(b) The conversion rate of UA pairs with ”deep” relationship.
Furthermore, we observe that the lifelong UA interactions have strong connection with the ultimate platform revenue. As shown in Fig. 2(a), the count of UA pairs with both follow and frequent gift (”deep”) relationship is positively correlated with the total revenue value. On the other hand, as users continue to interact with the authors with ”deep” relationship, its conversion rate will increase, as shown in Fig. 2(b). Hence, it provides insights to improve the platform profits by modeling the LIV of UA pairs.
To the best of our knowledge, this paper is the first to Reinforce Lifelong Interaction Value of User-Author pairs (RLIV-UA) using RL, based on each UA’s interaction and its corresponding cumulative reward. As shown in Fig. 1, the RLIV-UA model dynamically optimizes the LIV of an UA pair to improve the stickiness of both user U and author A progressively.
However, directly applying RL to model the LIV of UA pairs presents several challenges: First, compared with existing request-based RL applications in RS, the UA interactions may happen in requests with long time span. In practice, industrial RSs involve massive numbers of users and authors, resulting in an enormous and sparse UA state space, which limits the storage of all interaction traces of UA pairs over a long time period, such as three month.
@@
To solve the above problems, we propose a novel Sparse Cross-Request Interaction Markov Decision Process (SCRI-MDP) and introduce an Adjacent State Approximation (ASA) method to construct RL training samples. Moreover, we propose a Multi-Task Critic Learning (MTCL) (Liu et al., 2023b) architecture to model gradually stronger UA relationships, such as click → follow → gift. Finally, due to the sample sparsity of each UA pair and high variance of labels, an auxiliary supervised learning task is designed to improve the stability and convergence of the RLIV-UA model.
Overall, the main contributions of this paper are as follows:
  • A novel LIV model, i.e. RLIV-UA, is proposed to reinforce the lifelong interaction value of UA pairs for large-scale RS.
  • Based on SCRI-MDP, a novel ASA method is introduced to construct RL training samples.
  • An MTCL architecture is proposed to model not only trivial interactions like click and watch time, but also ”deep” interactions like follow and gift. To improve the convergence of the RLIV-UA model, an auxiliary supervised learning task is designed.
  • Results in offline and online experiments show that the RLIV-UA model can improve both user engagement and author benefits, thus improving platform profits.
  • The RLIV-UA has been successfully deployed on two short-video applications, i.e. Kuaishou and Kwai, which include over 400 million daily active users.

### v2

@@ Abstract.
Recommendation systems (RS) help users find interested content and connect authors with their target audience. Most research in RS tends to focus either on predicting users’ immediate feedback (like click-through rate) accurately or improving users’ long-term engagement. However, they ignore the influence for authors and the lifelong interaction value (LIV) of user-author pairs, which is particularly crucial for improving the prosperity of social community on different platforms. Currently, reinforcement learning (RL) can optimize long-term benefits and has been widely applied in RS. In this paper, we introduce RL to Reinforce Lifelong Interaction Value of User-Author pairs (RLIV-UA) based on each interaction of UA pairs. To address the long intervals between UA interactions and the large scale of the UA space, we propose a novel Sparse Cross-Request Interaction Markov Decision Process (SCRI-MDP) and introduce an Adjacent State Approximation (ASA) method to construct RL training samples. Additionally, we introduce Multi-Task Critic Learning (MTCL) to capture the progressive nature of UA interactions (click → follow → gift), where denser interaction signals are leveraged to compensate for the learning of sparse labels. Finally, an auxiliary supervised learning task is designed to enhance the convergence of the RLIV-UA model. In offline experiments and online A/B tests, the RLIV-UA model achieves both higher user satisfaction and higher platform profits than compared methods.
@@ 1. Introduction
The recommendation system (RS) aims to help users discover content aligned with their interests, while simultaneously enabling content authors to reach their target audiences, thereby facilitating fan accumulation and revenue generation (Zangerle and Bauer, 2022; Jesse and Jannach, 2021; Afsar et al., 2022). By fostering repeated, mutually beneficial interactions between users and authors, RS plays a pivotal role in cultivating vibrant platform ecosystems, ultimately driving increased user engagement, traffic, and commercial returns (Azaria et al., 2013; Dias et al., 2008; Lu et al., 2014; Ras et al., 2017; Jannach and Adomavicius, 2017; Jannach and Jugovac, 2019; Jannach and Bauer, 2020).
@@
Current research in RS largely falls into two categories. The first focuses on improving the accuracy of immediate user feedback prediction shown in Fig 1, such as click-through rate (CTR), at each recommendation request, typically using deep neural networks (DNNs) (Linden et al., 2003; He et al., 2014; Covington et al., 2016; Zhou et al., 2018; Gu et al., 2020). The second category leverages reinforcement learning (RL) to optimize long-term user engagement from the user’s perspective (Zhang et al., 2022; Xue et al., 2023; Chen et al., 2021; Wang et al., 2022; Cai et al., 2023; Zheng et al., 2018; Zou et al., 2019; Wang, 2021), dynamically maximizing session-level or trajectory-level cumulative rewards (Sutton, 2018; Van Hasselt et al., 2016; Fujimoto and Gu, 2021).
However, both paradigms largely overlook a critical dimension: the Lifelong Interaction Value (LIV) of user-author (UA) pairs. By neglecting the bidirectional, evolving nature of UA relationships, existing methods fail to capture how sustained interactions, from initial discovery to deep loyalty, jointly benefit both parties and, by extension, the platform itself. This omission is particularly consequential, as the long-term stickiness and vitality of platform communities hinge on nurturing these relationships.
As further evidenced in Appendix C, we observe a strong positive correlation between the depth and frequency of lifelong UA interactions and key platform-level outcomes, including total revenue and average app usage time. This underscores a fundamental insight: strategically modeling and reinforcing the LIV of UA pairs is not merely a user- or author-centric optimization problem, but a direct pathway to maximizing overall platform prosperity.
To the best of our knowledge, this paper presents the first framework to Reinforce Lifelong Interaction Value of User-Author pairs (RLIV-UA) using RL. As shown in Fig. 1, RLIV-UA dynamically optimizes the cumulative value of each UA interaction to progressively strengthen the mutual stickiness between user U and author A. Directly applying RL to model LIV, however, introduces significant challenges: (1) UA interactions often span long, irregular time intervals, unlike the dense, request-aligned interactions typically modeled in RS; (2) The combinatorial scale of the UA state space, driven by hundreds of millions of users and authors, renders it infeasible to store complete historical interaction traces over extended periods (e.g., months).
@@
To address these challenges, we propose a novel Sparse Cross-Request Interaction Markov Decision Process (SCRI-MDP) to formally model sparse, long-interval UA dynamics; an Adjacent State Approximation (ASA) method to construct practical RL training samples under storage constraints; and a Multi-Task Critic Learning (MTCL) architecture (Liu et al., 2023b) that captures the progressive nature of UA relationships (e.g., click → follow → gift), leveraging denser signals to bootstrap learning of sparse, high-value actions. Finally, to combat sample sparsity and label variance, we introduce a supervised learning task to stabilize training and accelerate convergence.
In summary, the key contributions of this work are:
  • We propose RLIV-UA, a novel reinforcement learning framework designed to optimize the lifelong interaction value of user-author pairs in large-scale recommendation systems.
  • We introduce SCRI-MDP, a formalism for modeling sparse, cross-request UA interactions, along with ASA, a practical method for constructing RL training samples under industrial constraints.
  • We design an MTCL architecture to model the hierarchical progression of UA relationships (e.g., from clicks to gifts), augmented by an auxiliary supervised learning task to ensure stable and efficient training.
  • Extensive offline simulations and large-scale online A/B tests demonstrate that RLIV-UA significantly improves both user engagement and author benefits, leading to substantial gains in platform revenue.
2. Problem Formulation
Existing RL methods in RS often model user behaviors as infinite request-level markov decision process (MDP) (Sutton, 2018). Specifically, the time interval between adjacent states $\Delta=1$ always holds. However, under the UA interaction space, the time interval between the same UA pair’s adjacent interactions satisfies $\Delta\geq 1$. For a specific UA pair, the interactions are usually sparse due to the large scale of candidate recommended items. Therefore, we define a novel sparse cross-request interaction MDP to model the LIV of UA pairs:
Definition 2.1 (Sparse Cross-Request Interaction MDP).
A Sparse Cross-Request Interaction MDP (SCRI-MDP) is a tuple represented as $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\Delta\rangle$, where:
  • State space $\mathcal{S}$: Each state $\mathbf{s}^{t}_{\mathrm{ua}}\in\mathcal{S}$ encodes user features $\mathbf{f}^{t}_{\mathrm{u}}$, author features $\mathbf{f}^{t}_{\mathrm{a}}$ and interaction features $\mathbf{f}^{t}_{\mathrm{ua}}$ (e.g., cumulative watch time, gift count).
  • Action space $\mathcal{A}$: The action $a^{t}_{\mathrm{ua}}\in\mathcal{A}$ is the weight of the ranking score at request $t$.
    (1) $final\_score=rank\_score*(c+a_{\mathrm{ua}}^{t}*w)^{b}$
    where $w,b,c$ are hype-parameters.
  • State transition distribution $\mathcal{P}$: $P(\mathbf{s}^{t+1}_{\mathrm{ua}}\mid\mathbf{s}^{t}_{\mathrm{ua}},a^{t}_{\mathrm{ua}},\Delta^{t}_{\mathrm{ua}})$ is determined by the time gap $\Delta^{t}_{\mathrm{ua}}\geq 1$ between consecutive interactions of the same UA pair, skipping intermediate requests for that pair if no interaction occurs.
  • Reward function $\mathcal{R}$: To model the progressive changes of UA relationship, several functions are designed to model the long-term reward of different immediate feedback between UA. Let $r_{\mathrm{ua},c},r_{\mathrm{ua},w},r_{\mathrm{ua},f},r_{\mathrm{ua},g}$ be the reward function of click, effective view, follow, and gift labels, respectively. $C=\{c,w,f,g\}$ denotes the target label set and $n=4$ denotes the cardinality of $C$.
    (2) $r_{\mathrm{ua},l}=\begin{cases}1,&\text{the behavior $l$ happens},l\in C\\ 0,&\text{otherwise}\end{cases}$
  • Discount factor $\gamma\in[0,1)$: Balances immediate and future rewards.
  • Interaction gap $\Delta^{t}_{\mathrm{ua}}$: Number of global requests between two consecutive interactions of the same UA pair.
Note that the proposed SCRI-MDP is a specific instantiation of the general semi-Markov Decision Process (semi-MDP) framework (Sutton et al., 1999). In a standard semi-MDP, actions can take variable amounts of time to complete, and state transitions occur upon the completion of these ’macro-actions’. In our case, the ’macro-action’ corresponds to the recommendation policy applied between two consecutive interactions of a UA pair. The decision point (state transition) only occurs when an actual UA interaction happens, which naturally fits the semi-MDP paradigm by handling the variable and often long time intervals ($\Delta\geq 1$) between interactions. Furthermore, the proposed SCRI-MDP can be regarded as an augmented MDP to convert the semi-MDP paradigm into a ’delay-free’ standard MDP paradigm. Then we introduces the following theorem.
Theorem 2.2 (Equivalence to augmented MDP).
Consider a Sparse Cross-Request Interaction MDP $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\Delta\rangle$ with interaction gap variable $\Delta^{t}_{\mathrm{ua}}\geq 1$. Define an augmented state $\bar{s}_{t}=(s^{t}_{\mathrm{ua}},\Delta^{t}_{\mathrm{ua}})$ where $\Delta^{t}_{\mathrm{ua}}$ is treated as part of the state vector, forming an augmented state space $\bar{\mathcal{S}}=\mathcal{S}\times\mathbb{N}$. Then $\mathcal{M}$ is equivalent to a standard MDP $\bar{\mathcal{M}}=\langle\bar{\mathcal{S}},\mathcal{A},\bar{\mathcal{P}},\mathcal{R},\gamma\rangle$, where the transition kernel is defined as:
(3) $\bar{P}((\mathbf{s}^{\prime},\Delta^{\prime})\mid(\mathbf{s},\Delta),a)=P(\mathbf{s}^{\prime}_{\mathrm{ua}},\Delta^{\prime}_{\mathrm{ua}}\mid\mathbf{s}_{\mathrm{ua}},\Delta_{\mathrm{ua}},a_{\mathrm{ua}}),$
and the reward function $\mathcal{R}$ remains identical. Any optimal policy $\pi^{*}$ in $\mathcal{M}$ corresponds to an optimal policy in $\bar{\mathcal{M}}$ under this mapping.
The proof of 2.2 is available in the Appendix A.1. Theorem 1 establishes a formal equivalence between the proposed SCRI-MDP and a standard MDP. It provides a rigorous justification for applying conventional MDP-based reinforcement learning algorithms (such as DQN, DDPG, or TD3, which are grounded in standard MDP theory) to solve the SCRI-MDP problem. By demonstrating that the SCRI-MDP can be transformed into an equivalent standard MDP through state augmentation, we ensure that the theoretical convergence properties and optimality guarantees of these well-studied algorithms remain applicable to our framework. This bridges the gap between our novel, problem-specific formulation and the vast body of existing RL theory and practice.
Based on the aforementioned SCRI-MDP formulation, we introduce a multi-task critic learning (MTCL) architecture to model the LIV of UA pairs by capturing the progressive changes of lifelong UA relationship. There are $n$ critic networks $Q_{\phi_{k}},k=1,...,n$ with $\phi_{k}$ as their trainable parameters to optimize the corresponding cumulative rewards $r_{\mathrm{ua},c},r_{\mathrm{ua},w},r_{\mathrm{ua},f},r_{\mathrm{ua},g}$. The final objective function is as follows:
(4) $\max_{\phi_{1},\cdots,\phi_{n}}\sum_{l\in C}\mathbb{E}\left[\sum_{t=0}^{\infty}{r_{\mathrm{ua},\,l}^{t}}\right]$
Each critic network is designed to learn from the states of UA pairs and corresponding rewards. Specifically, at one request of user $\mathrm{u}$, RS recommends author $\mathrm{a}$’s item by outputting the action $a^{t}_{\mathrm{ua}}\in\mathcal{A}$ which is the $t$-th interaction of $\mathrm{ua}$, then the optimal LIV of $\mathrm{ua}$ pair and action $a^{t}_{\mathrm{ua}}$ is defined as:
(5) $Q^{\ast}(\mathbf{s}^{t}_{\mathrm{ua}},a^{t}_{\mathrm{ua}})=r^{t}_{\mathrm{ua}}+\gamma\!\!\!\!\sum_{\mathbf{s}^{t+1}_{\mathrm{ua}}\in\mathcal{S}}\!\!\!\!P(\mathbf{s}^{t+1}_{\mathrm{ua}}\mid\mathbf{s}^{t}_{\mathrm{ua}},a^{t}_{\mathrm{ua}})\!\!\!\max_{a^{t+1}_{\mathrm{ua}}\in\mathcal{A}}\!\!Q^{\ast}(\mathbf{s}^{t+1}_{\mathrm{ua}},a^{t+1}_{\mathrm{ua}})$
3. Methodology
In this section, we propose the RLIV-UA model to represent the SCRI-MDP in the UA pair state space. Firstly, due to the long time span between adjacent states of UA pair in the SCRI-MDP, we propose the Adjacent State Approximation (ASA) method to construct RL training samples. Then, we introduce the detailed network architecture of multi-task LIV networks and the final online deployment of the RLIV-UA model. The algorithm pseudocode of the RLIV-UA model is presented in Appendix B.3.
3.1. Adjacent State Approximation
At the t-th interaction of a UA pair, the next state $\mathbf{s}^{t+1}_{\mathrm{ua}}$ is delayed until the next interaction which may occur much later. Thus, the traditional RL training sample $(\mathbf{s}^{t}_{\mathrm{ua}},a^{t}_{\mathrm{ua}},r^{t}_{\mathrm{ua}},\mathbf{s}^{t+1}_{\mathrm{ua}})$ cannot be formed in real time due to sparse, infrequent interactions. In industrial-scale recommendation systems, the sheer number of users and authors makes it infeasible to store the state of every UA pair in a key-value database. Instead, systems typically store user behavior as fixed-length sequential logs. Due to finite storage capacity, these sequences are truncated, which often prevents the system from retrieving a UA pair’s prior interaction, especially for infrequent or new relationships. In our online deployment, the success rate of constructing consecutive RL training samples $(\mathbf{s}^{t-1}_{\mathrm{ua}},a^{t-1}_{\mathrm{ua}},r^{t-1}_{\mathrm{ua}},\mathbf{s}^{t}_{\mathrm{ua}})$ by stitching historical logs is only 47%. Training LIV networks directly on such sparse and fragmented samples would lead to severe model bias, the RL agent learns from an incomplete, non-representative subset of UA trajectories, significantly degrading its ability to optimize long-term, bidirectional value. Moreover, learning from stitched historical tuples inherently delays the model’s ability to act on new relationships. Specifically, the very first interaction between a user and an author cannot be used for training until a second interaction occurs. This creates a critical blind spot: the model is unable to learn how to nurture the initiation phase of a UA relationship.
To address these challenges, we propose the Adjacent State Approximation (ASA) mechanism, a parameterized state predictor $f_{\theta}$ to reconstruct the missing next state from the current interaction:
(6) $\hat{\mathbf{s}}^{t+1}_{\mathrm{ua}}=f_{\theta}(\mathbf{s}^{t}_{\mathrm{ua}},a^{t}_{\mathrm{ua}},r^{t}_{\mathrm{ua}})$
where the state $\mathbf{s}^{t}_{\mathrm{ua}}=\{\mathbf{f}^{t}_{\mathrm{u}},\mathbf{f}^{t}_{\mathrm{a}},\mathbf{f}^{t}_{\mathrm{ua}},\Delta^{t}_{\mathrm{ua}}\}$ is augmented with the interaction gap $\Delta^{t}_{\mathrm{ua}}$, and $\theta$ denotes the trainable parameters of $f_{\theta}$.
Figure 2. Training sample construction for state predictor and RL model.
As illustrated in Fig. 2, we first train $f_{\theta}$ by retrieving the most recent interaction between the same UA pair from the user’s historical behavior logs. Industrial systems typically store interaction events (e.g., clicks, follows, gifts) as fixed-length sequences. To construct a supervised training sample, we first locate the timestamp $T$ of the user’s most recent interaction with author $a$. Since watching an item to completion typically marks the end of an interaction episode, we perform a binary search within a temporal window around $T$ to retrieve the corresponding action $a^{t-1}_{\mathrm{ua}}$ and associated feedback signals. This forms the training tuple $(\mathbf{s}^{t-1}_{\mathrm{ua}},a^{t-1}_{\mathrm{ua}},r^{t-1}_{\mathrm{ua}},\mathbf{s}^{t}_{\mathrm{ua}})$, optimized via mean squared error:
(7) $\mathcal{L}_{\theta}=\left\|f_{\theta}(\mathbf{s}^{t-1}_{\mathrm{ua}},a^{t-1}_{\mathrm{ua}},r^{t-1}_{\mathrm{ua}})-\mathbf{s}^{t}_{\mathrm{ua}}\right\|_{2}^{2}$
With ASA, then we can approximate the full RL training sample $(\mathbf{s}^{t}_{\mathrm{ua}},a^{t}_{\mathrm{ua}},r^{t}_{\mathrm{ua}},\hat{\mathbf{s}}^{t+1}_{\mathrm{ua}})$ for every interaction shown in Fig. 2, not just those with retrievable history. This directly addresses the 53% stitching failure rate and eliminates the cold-start blind spot for new UA pairs. Crucially, ASA enables the RL agent to learn from first-time interactions, capturing the full lifecycle of UA relationships, from initial discovery to deep, monetizable engagement, and ensuring comprehensive learning across the entire interaction space.
3.2. Theoretical Analysis
The proposed adjacent state approximation enables practical training of RL models in industrial recommendation systems. However, a critical question arises: How does the approximation error in ASA affect the optimality of the learned policy? To address this, we establish the following theorem that provides an error bound on the Q-function when using approximate state transitions.
Theorem 3.1 (Error Bound for Approximate State Transitions).
Let $Q^{*}$ be the optimal Q-function of the SCRI-MDP $\mathcal{M}$, and let $\hat{Q}$ be the Q-function learned using the ASA with maximum prediction error $\epsilon=\max_{s,a,r}\left|\hat{s}^{t+1}-s^{t+1}\right|$. If the ASA error $\epsilon$ is bounded and the reward function is Lipschitz continuous with constant $L_{r}$, then the difference between $\hat{Q}$ and $Q^{*}$ is bounded by:
(8) $\left|\hat{Q}(s,a)-Q^{*}(s,a)\right|\leq\frac{L_{r}\epsilon}{(1-\gamma)^{2}}$
for all $(s,a)\in\mathcal{S}\times\mathcal{A}$, where $\gamma\in[0,1)$ is the discount factor.
The proof of it is available in the Appendix A.2. This theorem provides a theoretical guarantee that the performance degradation caused by ASA is bounded and decreases as the ASA error $\epsilon$ decreases. In practice, we can train the ASA network $f_{\theta}$ to minimize $\epsilon$, ensuring that the learned policy remains close to optimal. Note that the Lipschitz continuity of the reward function is a standard and widely adopted assumption in theoretical RL literature to ensure the stability and boundedness of value functions. In our implementation, the reward functions are designed as bounded, non-negative functions of user engagement. This design choice inherently promotes smoothness in the reward landscape, making the Lipschitz assumption practically reasonable for our problem setting.
Figure 3. The training process of ASA.
We further validate the effectiveness of ASA by comparing it against a rule-based baseline, as illustrated in Fig. 3. The ASA method achieves a low MSE loss during training, which directly implies a small approximation error $\epsilon$ in estimating the next state representation $\mathbf{s}^{t+1}$. It further implies that the RL samples generated by ASA are of high fidelity. Consequently, when RLIV-UA is trained on these samples, it provably converges to a near-optimal policy, which satisfies the Q-value performance bound established in Theorem 3.1. More discussion and offline evaluation results of ASA is detailed in Appendix D.
Overall, ASA is a storage-constrained approximation mechanism designed for industrial-scale recommendation systems, where storing complete historical traces for all user-author pairs is infeasible. Unlike model-based RL approaches, e.g. Dreamer (Hafner et al., 2019), that learn transition dynamics for planning or sample efficiency, ASA serves a singular, pragmatic purpose: to reconstruct valid $(s,a,r,s^{\prime})$ tuples for RL training when the true next state $s^{\prime}$ is physically unavailable due to system limitations. Its objective is not to predict environmental dynamics, but to preserve training signal continuity under extreme data sparsity.